OpenStudy (anonymous):

can somone help me?? to prove tan3xtan2xtanx= tan3x-tan2x-tanx

OpenStudy (anonymous):

tan x + tan 2x + tan 3x - tan x tan 2x tan 3x = 0 => tan x + tan 2x + tan 3x ( 1 - tan x tan 2x ) = 0 => tan x + tan 2x = - tan 3x ( 1 - tan x tan 2x ) => ( tan x + tan 2x )/ ( 1 - tan x tan 2x ) = - tan 3x => tan 3x = - tan 3x => 2 tan 3x = 0 => tan 3x = 0 => 3x = 2kpi where k = 0, 1,2,3,4,5,.............. => x = 2 k pi/3 where k = 0 , 1, 2 ,3 , 4 , 5 , 6..............

OpenStudy (anonymous):

3x=2x+x applying tan b/s tan3x=tan(2x+x) tan3x=tan2x+tanx\1-tan2xtanx tan3x(1-tan2xtanx)=tan2x+tanx tan3x-tan3xtan2xtanx=tan2x+tanx tan3x-tan2x-tanx=tan3xtan2xtanx

OpenStudy (anonymous):

tysm!!!!!!